OpenStudy (anonymous):

Using complete sentences, explain how to factor each one. Be sure that the final factorization (or "answer") is a part of your explanation. 10x^2 - 7x - 3

OpenStudy (anonymous):

Help!

OpenStudy (anonymous):

let f(x)=10x^2 - 7x - 3 put x=1, f(1)=0 by factor theorem, x-1 is a factor of f(x) put x=-3/10, f(-3/10)=0 By factor theorm, 10x+3 is a factor of f(x) Therefore, f(x) can be factorized as (x-1)(10x+3)

OpenStudy (anonymous):

can u make it something looking like this : let me show you how. here we have ----> 2x^2 + 13x + 15 = so 'a' and 'c' here are 10 and 3 combined prime factors are 1,2,3,5 so let me group (1,2,3), and (5). i get (1x2x3)+(5) = 6 + 5 =11. this is not equal to 'b' that is 13. so we try another combination. lets try for (1,2,5) and (3). here, (1x2x5)+(3)= 13 so this one will work!! so we split 13x as 10x+3x so now i write the expression as: 2x^2+13x+15 = 2x^2 + (10)x + (3)x + 15 = 2x(x+5) + 3(x+5) = (2x+3)(x+5) <-------